

********************************************************************
UARS WEEKLY STATUS REPORT - 16 JAN 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
The half yearly calibration was successfully completed. In the coming
week, we plan to run field-of-view calibrations of SUSIM's lower
A poster paper entitled "Ultraviolet Irradiance at Solar Minimum" by
members of the SUSIM team was presented to the Washington, DC meeting
of the American Astronomical Society last week.
***********
PEM
***********
The PEM zenith spectrometers were powered January 4. HEPS-2 microprocessor
errors did not clear and the HEPS-2 unit was shut down. The HEPS-2 unit was
restarted on January 5 successfully. This type of problem has been seen before
and appears to be a symptom of being cold for a long period of time. It is not
anticipated that this problem will occur under solar array control. On
January 9, the PEM zenith instruments were shut down. During the time the PEM
zenith spectrometers are powered, PEM produced energy deposition profiles from
electron and proton precipitation.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS instrument has been powered off since Dec 27, 1997. MLS is scheduled to
begin north viewing operations on Jan 20, 1998.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING JAN. 16, 1998
WINDII is currently off, awaiting its next operating period
which is planned for Feb. 3 to 12.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 360 CD-ROM requests as of
16-January-98. There were five orders received and five orders shipped.
The CDHF previewed DAAC data for HALOE, HRDI, PEM, NMC, and UKMO
for the month of November 1997.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day SUSIM (6.3) Test (1R-1C=0L)
2 days SOLSTICE (4.52) Test (2R-2C=0L)
25 days PEM (3.32) Prod (1350R-775C=575L)
3 days MLS (4.90) Test (3R-3C=0L)
The CDHF continues to work with NSI to develop a support plan for the
entire UARSNET network, focusing on Instrument RACs first, redesign issues,
and then the global user community.
***********
FOT
***********
This report covers 03 January 1998 (Orbit 34492, GMT Day 003) through
09 January 1998.
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 004/1516 34516 ZEPS on
PEM 009/2327 34596 ZEPS off
The beta angle is decreasing toward a minimum angle of 0.0 degrees on
16 January 1998 (DOY 016).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently maintaining
greater than 25.0 V end-of-night (EON) load bus voltage. Battery temperatures
are stable with a temperature delta between Battery 2 and Battery 3 of 1.3
degrees C. Additional battery performance data for each battery on 03 January
1998 and 09 January 1998 are provided in Appendix D.
The clock error ranged from -10.9 to +10.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 01
% PEM 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 02
TRANSPONDER FREQUENCIES:
Transponder A: 97/364 - 2287.496935
Transponder B: 97/365 - 2287.499033
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 98/008 show the observatory data loss to be 20 hours, 41
minutes, 30.836 seconds (an increase of 06.144 seconds since last reported).
This is a 0.0375 percent data loss which equals a 99.9625 percent data capture
for the mission.
The increase of 6.144 seconds occurred over three days as follows:
7 a loss of 0.512 seconds on Day 005 due to a hit in the data,
7 a loss of 2.080 seconds on Day 006 due to a hit in the data,
7 a loss of 2.944 seconds on Day 007 due to insufficient overlap and a hit
in the data, and
7 a loss of 0.608 seconds on Day 008 due to a hit in the data.
No testing of the new CMS software release (R7B1V2) occurred this week.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer systems.
Production level-0 file transfers between CDHF and UMPGAL failed partially on
07 January 1998 at approximately 02:00 UTC. This was the only file transfer
problem identified this week.
The UARS simulator's Q-channel line problem was corrected on 05 January 1998.
Arrangements were made for NASCOM hardware support to be available to
troubleshoot the problem. After an initial failure on 05 January, NASCOM
support made corrections and successful Q-line channel dumps were obtained.
Continuous successful dumps were obtained throughout the day on Monday and
again on Tuesday. The corrective action taken by NASCOM to correct this
problem is not clear.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software has been
initiated. This week has been focused on the redesign (i.e., PDL); however,
coding is not expected to begin until next week. The coding should be done
within two weeks after that, provided no problems are encountered. Another
week or so is anticipated in order to complete database updates and testing.
Uplink scheduling then becomes contingent upon the Applications Processor (AP)
software changes to support the reduced-size Absolute Time Command Processor
(ATCP). In summary, the software will be ready for uplink by late February
provided no unanticipated problems are encountered.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
All of the 1997 Anomaly Reports have been processed to closure.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 January 1998 at
016/13:25:18 UTC.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 January 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34506 003/2328 98-002 HRDI Power OFF
34507 004/0043 98-003 PEM Power ON
34516 004/1516 98-004 PEM Power ON HEPS 2
34517 004/1650 98-005 OBC Clk Adj fr 20.6 to 0.00 msec/day
34550 006/2158 98-006 OBC Clk Adj fr 0.00 to 20.6 msec/day
34576 008/1530 98-007 OBC Clk Adj fr 20.6 to 0.00 msec/day
34596 009/2327 98-009 PEM/ZEPS Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 January 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTMP Y-HI HRDI IFE SURVIVAL HTR TEMP 34508
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 34579,93-6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
(none this report period)
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 January 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34517 004/16:54 (04 Jan 98) from 20.6 to 00.0 msec/day
34550 006/22:04 (06 Jan 98) from 00.0 to 20.6 msec/day
34576 008/15:33 (08 Jan 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
January 1998
03 Jan 1998 - GMT Day 003 Beta = 33.4 deg, SA Pos = 269 deg
E
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- --
------ ------- ------- ------
1 +728.0/+728.0 -2.73/-3.99 +0.0/ +0.0 18.6 0.000 0.0
2 +11.2/ -50.4 +6.44/+5.15 +31.2/-16.0 25.1 1.026 22.4
3 +33.6/ -22.4 +5.15/+3.87 +29.6/-16.0 25.0 1.026 21.7
09 Jan 1998 - GMT Day 009 Beta = 22.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +5.6/ -67.2 +6.44/+5.15 +32.4/-15.2 25.0 1.027 22.8
3 +67.2/ -11.2 +5.15/+3.87 +31.2/-16.0 24.8 1.023 22.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 January 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-001 FOT 34517 004/1650 14170 Generic Late Acquisition
98-003 FOT 34543 006/1006 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-124 FOT (EIGHT) DEC 97 14170 DEC. GEN LATE ACQ SUM
97-127 FOT 34207 349/00:18 GYRO GLITCH
97-128 FOT 34442 364/16:28 TAC CRASH
97-129 FOT 34444 364/20:03 TAC CRASH
97-130 FOT 34457 365/16:10 19647 TAC CRASH
(END 1997 AR DATA - ALL 1997 AR CLOSED)
98-002 FOT 34465 001/07:14 TAC 2 CRASH
98-003 FOT 34563 006/10:06 TAC 4 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period